A beam of light converges at a point P.Now a lens is placed in the path of the convergent beam 12 cm from P. At what point does the beam converge, if the lens is (i)A convex lens of focal length 20cm. (ii) A concave lens of focal length=16cm ?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Nikita
Object distance, u = +12 cm (virtual object)
(a) Focal length of the convex lens, f = 20 cm
     Image distance = v
     Applying lens formula
     1/v – 1/u = 1/f
     Therefore putting the values and solving the equation, we get v = 60/8 i.e. = 7.5 cm
     Hence, the image is formed 7.5 cm away from the lens, toward its right.
(b) Focal length of the concave lens, f = -16
     again image distance = v
     the again as previous, applying the lens formula, substituing the values of u and f, we get v= 48 cm.
     Hence, the image is formed 48 cm away from the lens, towards its right.
Regards
Arun (askIITians forum expert)